the length of a rectangle is twice its width if the length decreases by 6cm and the width increases by 5 cm the rectangle becomes a square find the area of the rectangle

Answer by checkley79(3341)

L=2W
(L-2)=(W+5)
(2W-2)=(W+5
2W-W=5+2
W=7 ANS.
L=2*7 ANS FOR THE RECTANGLE.
L=14 ANS. FOR THE RECTANGLE.
PROOF:
14-2=7+5
12=12 MAKES IT A SQUARE.
AREA=7*14=98 cm. IS THE AREA OF THE RECTANGLE.